OpenStudy (anonymous):

Simplify: log x + log(x-3) = 1

OpenStudy (anonymous):

Sorry teh question was to solve for x

OpenStudy (anonymous):

\[\log(x(x-3))=1\] is a start

myininaya (myininaya):

here's a start \[\log|(x(x-3))|=\log(10)\]

OpenStudy (anonymous):

then write in equivalent exponential form and write \[x(x-3)=10\] etc

OpenStudy (anonymous):

couldn't help yourself could you?

myininaya (myininaya):

i need to go to sleep for real

OpenStudy (anonymous):

simplify to log x^2-3x = 1 automatic base 10, so 10=x^2-3x so \[x ^{2}-3x-10=0\] solve for x (x-5)(x+2)=0 x=5, -2

OpenStudy (anonymous):

yeah those early morning anti derivatives come at you hard

myininaya (myininaya):

-2 won't work

myininaya (myininaya):

just x=5

OpenStudy (anonymous):

and don't forget to discard -2 because you cannot take the log of a negative number

OpenStudy (anonymous):

what myininaya said

OpenStudy (anonymous):

yes, that's true. forgot about that, sorry

myininaya (myininaya):

satellite when i put my mouse clicker over your picture it makes it look like your bicycle is going

myininaya (myininaya):

math help you have any questions?

OpenStudy (anonymous):

how did you start off with logx^2-3x = 1...what log rule is taht

myininaya (myininaya):

we have log(x)+log(x-3) right? on the left hand side? we can use ln|x(x-3)| but that equals ln|x(x)-x(3)|=ln|x^2-3x|

myininaya (myininaya):

see i use the distributive property inside the ln function

OpenStudy (anonymous):

oh i see..tahts makes sense now.....Can one of you help me with another solve for x problem?